Use the table provided in the handbook. Marsh Terban bought a home for $120,000 with a down payment of $30,000. Her rate of interest is 12 1/2 percent for 35 years. Calculate her (A) monthly payment, (B) first payment broken down into interest and principal and (C) balance of mortgage at end of month.
Answer #1

Home Cost = $ 120000, Down Payment = $ 30000, Mortgage = 120000 - 30000 = $ 90000

Tenure = 35 years or (35 x 12) = 420 months and Interest Rate = 12.5 %

Applicable Monthly Interest Rate = 12.5/12 = 1.04167 %

Let the monthly mortgage payments be $ m

Therefore, 90000 = m x (1/0.0104167) x [1-{1/(1.0104167)^(420)}]

90000 = m x 94.7639

m = 90000 / 94.7639 = $ 949.729 ~ $ 949.73

Interest Accrued over 1st Month = Beginning Mortgage Balance x Monthly Interest Rate = 90000 x 0.0104167 = $ 937.5

Interest Portion of 1st Monthly Payment = Interest Accrued over 1st Month = $ 937.5

Principal Repaid = 949.729 - 937.5 = $ 12.23

Balance of Mortgage post 1st Month = 90000 - 12.23 = $ 89987.77